OpenStudy (anonymous):

I dont understand how to factor this problem. 16s^4 - 81. Can somebody show me and explain this to me?

OpenStudy (anonymous):

whats the biggest number that goes into both 16 and 81?

OpenStudy (anonymous):

Hint 3^4 = 81

OpenStudy (anonymous):

you could use a difference of 2 squares

OpenStudy (anonymous):

This is the difference of two terms to the fourth power. It is of the form x^4 - y^4 which is factored (x - y)[x^3 + (x^2)(y) + (x)(y^2) + y^3] Here, "x" is 2s and "y" is 3.

OpenStudy (anonymous):

The second factor, [x^3 + (x^2)(y) + (x)(y^2) + y^3], can itself be factored into (x + y)(x^2 + y^2). So, you should get 3 factors when you are done.

OpenStudy (anonymous):

Whats the biggest number between 16 and 81?

OpenStudy (anonymous):

Is your last question related to your question on factoring? In any case 16 = 2^4 and 81 = 3^4. Perhaps you are looking for "biggest number to the fourth power" in an interval?